Samantha Jones has a job as a pharmacist earning $30,000 per year, and she is deciding whether to take another job as the manager of another pharmacy for $40,000 per year or to purchase a pharmacy that generates revenue of $200,000 per year. To purchase the pharmacy, Samantha would have to use her $20,000 savings and borrow another $80,000 at an interest rate of 10 percent per year. The pharmacy that Samantha is contemplating purchasing has additional expenses of $80,000 for supplies, $40,000 for hired help, $10,000 for rent, and $5,000 for utilities. Assume that income and business taxes are zero and that the repayment of the principal of the loan does not start before three years.(b) Suppose that Samantha expects that another pharmacy will open nearby at the end of three years and that this will drive the economic profit of the pharmacy to zero. What would the revenue of the pharmacy be in three years?(d) Suppose that Samantha expects to see all the pharmacy at the end of three years for $50000 less that the price she paid for it and that she requires a 15 precent return on her investment. Should she still purchase the pharmacy?
1
Expert's answer
2013-03-01T04:42:07-0500
(a) Economic profit for year will be: 200,000 - (80,000 +40,000 + 10,000 + 5,000 + 8,000 + 40,000) = 17,000, so it will be profitable,
but not better, than to work as a manager.
(b) The revenue will be 200,000 - 17,000 = 183,000
(d) It will be still better to work as a manager, because all the costs are higher than 40,000 salary.